Your uncle has a life expectancy of another 15 years. He wants $25,000 per year to supplement his social security payments to live on. If interest rates are 4%, how much money must he have now to get the $25,000 per year for the next 15 years?
Answer #1

Answer : In 15 years the total amount becomes,

Amount (A) = $25,000 * 15 = $375,000

Time (t) = 15 years

rate of interest (r) = 4% = 0.04

Now, we know that for annual interest rate,

A = P * (1 + r)^t

=> 375,000 = P * (1 + 0.04)^15

=> 375,000 = P * 1.8

=> P = 375,000 / 1.8

=> P = 208,333.33

=> P = 208,333 (approximately).

Therefore, my uncle have to $208,333 (approximately) money at present situation to get $25,000 per year.